OpenStudy (anonymous):

Two trains leave the station at t=0 and travel with constant velocity V along straight tracks that make an angle Theta. a) show that the trains are separating from each other at a rate of v(sqrt(2-2cosTheta)) b)What does this formula give for Theta=pi?

OpenStudy (anonymous):

v=constant a=0

OpenStudy (anonymous):

\[x=vt\] for both cars

OpenStudy (anonymous):

angle between both cars 180-theta

OpenStudy (anonymous):

\[r^2=2(vt)^2-2(vt )^2\cos (\pi -\theta)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

when v= pi cos pi=-1 which gives 2v as the result

OpenStudy (anonymous):

\[r=\sqrt{(vt)^2(2+2\cos \theta )}= vt \sqrt{2+2\cos \theta} \] \[dr/dt=v=r/t\]

OpenStudy (anonymous):

\[v=\frac{ vt \sqrt{2+2\cos \theta}}{ t }=v \sqrt{2+2\cos \theta}\]

OpenStudy (anonymous):

\[v \sqrt{2+2\cos \pi}=v \sqrt{2+2(-1)}=0\]